OpenStudy (anonymous):

which of the following equation(s) describes these points (CHOOSE MULTIPLE ANSWERS CHOICES) (-6,4) & (3,1) A. y - 4 = -1/3(x+6) B. y - 1= -3(x-3) C. y - 3 = -1/3(x-1) D. y - 1= -1/3(x-3) E. y - 4 = -1/3(x-6) F. y - 4 = -3(x+6)

hartnn (hartnn):

do u know how to find slope of lines given 2 points ?

OpenStudy (anonymous):

yes... x1, y1, x2, y2

hartnn (hartnn):

so what slope u got ?

OpenStudy (anonymous):

y = -0.3333333333333333x + 2

hartnn (hartnn):

thats correct! so u got slope = m =-1/3

hartnn (hartnn):

u have point 3,1 use, y-y1 = m(x-x1) put x1 = 3,y1 =1, m=-1/3 what u get ?

OpenStudy (anonymous):

lol omg what i read sound so confusing...im sorry, im not really good in math !

hartnn (hartnn):

so which option u got as correct answer ?

OpenStudy (anonymous):

B. y - 1= -3(x-3) & C. y - 3 = -1/3(x-1)

hartnn (hartnn):

no.. y-1 = (-1/3 )(x-3) its D.